OpenStudy (anonymous):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (anonymous):

\[\frac{1}{3} \left(\frac{3}{5}+\frac{3}{4}\right)=\frac{9}{20} \]